Please Help:
Find the compound interest earned by the deposit. Round to the nearest cent.
$11,000 at 11% compounded annually for 16 years
A) $18,150.00
B) $47,419.84
C) $41,630.48
D) $19,360.00

Answer by rfer(16322)

A=P*(1+(r/n))^(1*16)
A=11000*(1+.11)^16
A=11000*5.310894=58419.84-11000=47419.84 (B)